#6 Two years ago, the price of a bond was 1,083.45 dollars, and one year ago, the price of the bond was 1,146.68 dollars. Over the past year, the bond paid a total of 66.2 dollars in coupon payments, which were just paid. If the bond is currently priced at 1,128.06 dollars then what was the rate of return for the bond over the past year (from 1 year ago to today)? The par value of the bond is 1,000 dollars. Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.

#7 Bonds issued by Fairfax Mechanical were priced at 906.31 dollars six months ago and are priced at 898.23 dollars today. The bonds have a face value of 1,000 dollars, pay semi-annual coupons, and just made a coupon payment. The bonds had a percentage return over the past six months (from 6 months ago to today) of 7.02 percent. What is the coupon rate of the bonds? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.

hand & calculator plzz ThankYou
Answer #1

6: Rate of return on bond = (Price at end-Price at the beginning+ Coupon)/Price at the beginning

= (1128.06-1146.68+66.2)/ 1146.68

=4.15%

7:

Rate of return on bond = (Price at end-Price at the beginning+ Coupon)/Price at the beginning

7.02% = (898.23-906.31+Coupon)/ 906.31

0.0702 = (-8.08+ Coupon)/ 906.31

(-8.08+ Coupon)= 906.31*0.0702

(-8.08+ Coupon)= 63.62

Coupon =71.70

Annual coupon = 71.7*2 = 143.41

Hence coupon rate = 14.34%